OpenStudy (zolock):

Show all your work. 1. Refer to the triangle below. If a = 3 and b = 3. a. Find the exact length of c in unsimplified form b. Find the exact length of c in simplified radical form. c. Find c to the nearest hundredth. ( please help I need to turn it in)

OpenStudy (ipwnbunnies):

Is it a right triangle?

OpenStudy (zolock):

yes

OpenStudy (anonymous):

the Pythagorean theorem right? a^2+b^2=c^2 so a^2+b^2=c^2 3^2+3^2=c^2 9+9=c^2 18=c^2 c=square root of 18. 1a) c=square root of 18. <--unsimplified form 1b) square root of 18 =3 square root of 2. <--simplified radical form 1c) square root of 18 =4.24264...

OpenStudy (anonymous):

apply teh pytahgors theorem simple